OpenStudy (anonymous):

A silicon chip used in an integrated circuit of a computer has a mass of 4.86 mg. How many silicon atoms are present in this chip?

OpenStudy (anonymous):

@ganeshie8 Please help me urgent !

OpenStudy (julian.pitalua):

28.08 g /6.02E23 atoms =0,00486g/ xxx atoms

OpenStudy (anonymous):

Can I solve it like this ? no. of moles of element = no. of particles / 6.22 * 10^23 particles No. of moles of a element : mass of element in g / molar mass of element in g/mol = 4.86 / 28 = 0.1736 * 6.22 * 10^23 = 1.079792e23

OpenStudy (anonymous):

@julian.pitalua Can you give one more reply here !

OpenStudy (anonymous):

@ganeshie8 Can you help me solve the problem please ?

OpenStudy (julian.pitalua):

how did you get 0.1736? Yes, you can do it like that, but it would be this way: 4.86 / 28 = X /6.22 * 10^23